A quantity of gas occupies a volume of 0.4M² a pressure of 100KN/M³ and a temperature of 20°C the gas is compressed isothermally to a pressure of 450KN/M² then expanded adiabatically to its intial volume for this quantity of gas determine;

1:the heat transferred during the compression

2:the change of internal energy during the expansion

3:the mass of gas.


1
Expert's answer
2021-05-31T06:08:50-0400

"n=\frac{R_1V_1}{RT_1}=\frac{100*10^3*0.4}{8.314*293}=16.42mol"

"v_2=\frac{nRT}{P_2}=\frac{16.4*8.314*293}{450*10^3}=0.08878m^3"

(1) "Q_c=nRT_1\ln(\frac{V_2}{V_1})=16.42*8.314*293\ln(\frac{0.08878}{0.4})"

"=-60210.85kJ"

(2)"\triangle u=\frac{nR(T_3-T_2)}{y-1}"

But "V_3=V_1=0.4"

"P_3=P_2(\frac{V_2}{V_3})^y=450*10^3(\frac{0.8878}{0.4})^{1.4}"

"=54697.78"

"T_3=9.293(\frac{54697*78}{450*10^3})^{\frac{0.4}{1.4}}=160.46"

"\triangle u=\frac{16.42*8.314(160-293)}{1.4-1}=-45391.53kJ"

(3) "PV=mRT=m=\frac{100*0.4}{8.314*293}=0.0164kg"